If Catman is around please do this one, everybody is giving me wrong answers except you. It is not 21.875 nor 21.876.
Homework: HW 6 (Ch 10 & 13) Save Score: 0.25 of 1 pt 19 of 21 (18 complete) Hw Score: 75.99%, 15.96 of 21 p 13.2.27 EQuestion Help A certain type of candy is reported to be distributed with the color percentages shown in the table below. The number of each color in a random bag of candy is also shown below. Use the chi-square goodress-of-fit test to decide whether the color distribution differs from the reported distribution. Use α-0.05. Color Frequency Green 15 17 Orange 21 26 Blue 24 25 Red 13 14 Brown 13 14 Ha: The frequency is different for each color yB. Ho: The color distribution is the same as that reported Ha: The color distribution is different from that reported Ho: The color distribution is different from that reported Ha: The color distribution is the same as that reported C. D. Ho: The frequency is different for each color Ha: The frequency is the same for each color Compute the value of the test statistic, 2 2-21 876 (Round to three decimal places as needed.) Enter your answer in the answer box and then click Check Answer. 2 remainina Clear All Final Check ec the following data on the number of acute postoperative days in the hospital using the dynamic
Answer #1

Ans:

Total number of observations=17+26+8+25+14+6=96

Expected count=pi*96

(pi is the percentage for respective colur)

Observed(O) Expected(E) (O-E)^2/E
Green 17 14.4 0.469
Orange 26 20.16 1.692
Yellow 8 13.44 2.202
Blue 25 23.04 0.167
red 14 12.48 0.185
Brown 6 12.48 3.365
Total= 96 96 8.080

Chi square score=8.080

df=6-1=5

p-value=CHIDIST(8.080,5)=0.1519

As,p-value>0.05,we fail to reject H0.The color distribution is the same as reported.